solve the following. one more than5,two more than6,one less than5,two less7
Found 2 solutions by CubeyThePenguin, ikleyn:
Answer by CubeyThePenguin(3113)

one more than 5 = 5 + 1 = 6

two more than 6 = 6 + 2 = 8

one less than 5 = 5 - 1 = 4

two less than 7 = 7 - 2 = 5

Answer by ikleyn(52781)

.

For such assignments, use your fingers.